OpenStudy (anonymous):

On the graph of the equation 3x + 2y = 18, what is the value of the y-intercept?

OpenStudy (anonymous):

y-intercept::put x=0

OpenStudy (anonymous):

3(0)+2y=18,y=9

OpenStudy (anonymous):

3x+2y=18 Write the equation 2y=-3x+18 Subtract 3x to get -3x y=-3/2x+9 Divide each term by 2 to isolate the y. The form for slope intercept form is y=mx+b m= slope and b= y-intercept b=9

OpenStudy (whpalmer4):

If eq is not already in slope intercept form, setting x=0 and solving for y is usually a quick route to the y intercept. Set y=0 and solve for x to find x intercept. If eq in standard form like this one, \[Ax +By=C\] Y intercept is C/B X intercept is C/A But you have to remember which is which.